Grade 12th passPhysical Chemistry

Find the volume of CO2 at NTP obtained on heating 10 g of 90% pure limestone (1) 22.4 L(2) 20.16 L(3) 2.24 L(4) 2.016

Vikash Kumar Singh
8 Years agoGrade 12th pass

1 Answer

Neeraj Kumar
8 Years ago
90% pure limestone means it have 90gm limestone in 100gm sample.this means that we have 9gm limestone in 10gm sample.So moles of lime stone is=9/100=0.09molesNow,CaCO3--> CaO + CO21mole CaCO3 gives 22.4L CO20.09moles CaCO3 gives = 22.4×0.09=2.016LHence, the answer is 2.016L